Greenhab Report – December 10th

Crew 217 GreenHab Report – 10-12-2019

GreenHab Officer: Cynthia Hori

Environmental control:

Ambient with door open

Heating

Shade cloth (40%) on

Average temperatures

Low 65 F

High 90 F

Hours of supplemental light :5

Daily water usage for crops: 80 L

Daily water usage for research and/or other purposes:

Water in Blue Tank— 300 gallons

Time(s) of watering for crops: 5 pm

Changes to crops: Pepper seedlings emerging

Narrative: A second shade cloth (30%) will be added tomorrow and squash will be transplanted. New lettuce seeds going in tomorrow.

Harvest: none

Operations Report – December 10th

Crew 217 Operations Report 12/10/2019

SOL: 1

Name of person filing report: Hope Zubrin

Non-nominal systems: radios

Notes on non-nominal systems: batteries not charging

Generator:

Hours run: 12

From what time last night: 7:00 pm

To what time this morning: 7:00AM

List any additional daytime hours when the generator was run: n/a

Solar— SOC 70 % last night (Before generator is run), 72% tonight

Diesel Reading – not checked

Station Propane Reading –not checked

Ethanol Free Gasoline: 5 gallons

Water (loft tank): 25 gallons

Water Meter: 1467244 units

Water (static tank): 425 gallons

Static to Loft Pump used – yes

Water in GreenHab: 300 gallons

Water in ScienceDome: zero gallons

Toilet tank emptied: no

Deimos rover used: off campus

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:yes

Hours: 116.6

Beginning charge: 100 %

Ending charge: 77%

Currently charging: Yes

Opportunity rover used: off campus

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours:121.6

Beginning charge:100%

Ending charge:63%

Currently charging: yes

Notes on rovers: Curiosity ignition switch tightened as it had become loose

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:not used

Oil Added? n/a

ATV Fuel Used: zero gallons

# Hours the ATVs were used today:

Notes on ATVs: nothing to report

HabCar used and why, where yes to town for supplies

CrewCar used and why, where?In Grand Junction

General notes and comments:

Summary of internet: troubles downloading

Summary of suits and radios: all suits operational see radio notes

Summary of Hab operations: nothing to report

Summary of GreenHab operations: we adjusted the grow lights and are working on temperature moderation

Summary of ScienceDome operations: replaced table

Summary of RAM operations: nothing to report

Summary of any observatory issues: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: nothing at this time.

Greenhab Report – December 9th

Shannon Rupert

8:44 PM (30 minutes ago)
to MDRS
Crew 217 GreenHab Report – 09-12-2019

GreenHab Officer: Cynthia Hori

Environmental control:

Ambient with door open

Heating

Shade cloth (40%) on

Average temperatures: 20 degrees C

Low

High

Hours of supplemental light :5

Daily water usage for crops: 80 L

Daily water usage for research and/or other purposes:

Water in Blue Tank— 300 gallons

Time(s) of watering for crops: 5 pm

Changes to crops: Pepper seedlings emerging

Narrative: A second shade cloth (30%) will be added tomorrow

Harvest: Radishes were thinned and 141 g removed and eaten at dinner–

Sol Summary – December 10th

Crew 217 Sol Summary Report 10-DEC-2019

Sol: 1

Summary Title: The First Day on Mars aka Taco Tuesday

Author’s name: Shannon Rupert

Mission Status: Our Mission is in full swing. We awoke to our first Martian day and made the most of it.

Sol Activity Summary: We spent the morning sharing our teaching experiences and activities, and at one point while I was smashing #10 cans on the upper deck, the crew was busy downstairs creating a habitat out of cups, plates, straws, file folders and recycled plastic containers. I was reminded of that Apollo 13 scene where they dump everything on the table to create a solution for the scrubber problem (“Failure is not an option”). After lunch, we trained in suits and rovers while taking our first short EVAs to the Marble Ritual. Everyone was super excited about wearing the suits and being outside. Then we settled down to a wonderful Taco Tuesday dinner with homemade tortillas and space ball taco meat. We created our work plan for tomorrow, then decided to spend the rest of the evening doing COMMS and relaxing. It’s been a wonderful but busy day.

Look Ahead Plan: We will spend the morning in the lab making plates for our halophile experiment. Then part of the crew will go on an EVA to Lith Canyon to collect pothole water samples for analysis and they will also collect samples from a green layer in the stratigraphy here and we will plate them to see what is living there. Finally, a second EVA will go to The Moons to collect gypsum samples to plate.

Anomalies in work: Atila and I needed to go to town for supplies and that set us back a couple of hours so that is why we pushed the plate making back until morning. It’s hard to create a Martian pace with only three days to do so much, but I am trying.

Weather: Sunny and very warm for this time of year

Crew Physical Status: Happy, healthy, some of us are clean, all of us will sleep well.

EVA: Two short training EVAs to the Marble Ritual, XO lead both of them.

Reports to be filed: Sol Summary, Operations Report. GreenHab Report, Operations Report, HSO Beginning of SIM checklist, Bios/Photos, Mission Plan, Journalist Report and Photos.

Support Requested: Teachers need to be added to the eList. We may not be able to send all reports this evening as we don’t know how much internet we used this morning. We have decided to turn off the wifi from 6 am to 7 pm every day as everyone still wants to use their phones for photos and in this way, we won’t have to worry about using the internet.

Journalist Report – December 10th

Crew 217 Journalist Report
10-Dec-2019

Teresa Hislop

Sol Two

First day of full simulation. SUCCESS! No deaths, infinite
enthusiasm. One could say our experiences were out of this world.

We spent the morning sharing ideas, lesson plans, and educational philosophies. Everyone’s approach to Mars education differed but none were alien. Kevin’s students create to scale Mars habitats, Jen uses Mars to hook ELA readers, and Hope takes students on a Mission to Mars for an entire semester. Ann’s Space Colonization curriculum is on her website and Cynthia shared NASA and Teacher-Pay-Teacher resources. Unquestionably, one of the most valuable NSBU outcomes will be the networking and sharing of resources and ideas that occurred this morning.

The afternoon took us outside for our first EVA (extra vehicular activity). In two sets of four, we suited up, drove the Rovers, and explored the Martian landscape. Except for a minor delay when the first crew had to return from airlock to correct a communications glitch, the EVAs were without incident.

The Internet continues to be the most non-stellar aspect of the experience. Yesterday it took 15 attempts to send the journalist report. None of the other officers were able to submit reports. Today connectivity continues to be spotty and bandwidth thin.

Kevin and Jen created an amazing mission patch. Jen, as Health and Safety Officer, has checked every fire alarm and carbon monoxide monitor in the complex. Ann completed tomorrow’s EVA requests. Kevin created and completed a science report. Yesterday Cynthia watered plants and planted lettuce. Teresa thinned the radishes and Ann used the radish greens in the evening’s stir fry. Tomorrow’s EVA adventures will take us gypsum hunting, pot hole sampling, and green stone collecting.

Greenhab Report – December 10th

Shannon Rupert

8:44 PM (30 minutes ago)
to MDRS
Crew 217 GreenHab Report – 09-12-2019

GreenHab Officer: Cynthia Hori

Environmental control:

Ambient with door open

Heating

Shade cloth (40%) on

Average temperatures: 20 degrees C

Low

High

Hours of supplemental light :5

Daily water usage for crops: 80 L

Daily water usage for research and/or other purposes:

Water in Blue Tank— 300 gallons

Time(s) of watering for crops: 5 pm

Changes to crops: Pepper seedlings emerging

Narrative: A second shade cloth (30%) will be added tomorrow

Harvest: Radishes were thinned and 141 g removed and eaten at dinner–